OpenStudy (anonymous):

Find the proportionality constant (or constant of variation) Q. V is directly proportional to the cube of R, and V=81 when R=3.

OpenStudy (unklerhaukus):

\[V\propto\sqrt[3]{R}\\V=c\times\sqrt[3]{R}\]